### v2

# On the Number of Degenerate Simplex Pivots

Abstract
The simplex algorithm is one of the most popular algorithms to solve linear programs (LPs). Starting at an extreme point solution of an LP, it performs a sequence of basis exchanges (called pivots) that allows one to move to a better extreme point along an improving edge-direction of the underlying polyhedron.
A key issue in the simplex algorithm’s performance is degeneracy, which may lead to a (potentially long) sequence of basis exchanges which do not change the current extreme point solution.
In this paper, we prove that it is always possible to limit the number of consecutive degenerate pivots that the simplex algorithm performs to $n-m-1$, where $n$ is the number of variables and $m$ is the number of equality constraints of a given LP in standard equality form.
1 Introduction
The simplex algorithm first introduced in [12] has always been one of the most widely used algorithms for solving linear programming problems (LPs). Although it performs as a linear time algorithm for many real world linear programming models (see, e.g., the survey [33]) and a variant of it (the shadow simplex) has been shown to have polynomial smoothed complexity (see, e.g., [7, 8, 10, 24, 34, 39]), no polynomial version of the simplex algorithm has been found until these days, despite decades of research.
The simplex algorithm relies on the concept of basis, where a basis corresponds to an inclusion-wise minimal set of tight constraints which defines an extreme point solution of the underlying linear program. In each step, it performs a basis exchange by replacing one constraint currently in the basis with a different one. Such an exchange is called a pivot. From a geometric perspective, a basis exchange identifies a direction to move from the current extreme point. The simplex algorithm only considers pivots that yield an improving direction, with respect to the objective function to be optimized. As the result of pivoting, it is possible for the algorithm to either stay at the same extreme point, or to move to an adjacent extreme point of the underlying polyhedron. We refer to a pivot of the former type as degenerate, in contrast to a pivot of the latter type, which will be called non-degenerate. The pivot rule which determines how to perform these basis exchanges (i.e., deciding which inequality enters and which inequality leaves the basis) is the core of the simplex algorithm. Various pivot rules have been introduced over decades, but none of them yields a polynomial time version of the simplex algorithm so far (see, e.g., [3, 13, 14, 15, 17, 35, 21, 40] and the references therein).
When analyzing the performance of the simplex algorithm, one can often assume that the underlying polyhedron $\mathcal{P}$ is non degenerate, e.g., by performing a small perturbation of the constraints right-hand-side. In this case, a necessary condition to guarantee an efficient performance of the simplex algorithm, is that there is always a way to reach an optimal solution within a polynomial number of non-degenerate pivots. Geometrically, this would imply that there is always a path between two extreme points on the 1-skeleton of $\mathcal{P}$ (the graph where vertices correspond to the extreme points of $\mathcal{P}$, and edges correspond to the 1-dimensional faces of $\mathcal{P}$), whose length is bounded by a polynomial in the input description of $\mathcal{P}$. This is a major and still open question in the field (see for example [6, 11, 23, 31, 37]).
On the other hand, even in the case when a short path exists, it is not clear how the simplex algorithm would be bound to follow it (see for instance the case of $0/1$-polytopes discussed in [4]). In fact, given any LP one can apply some transformation to it to ensure the existence of a short path between an initial extreme point solution and an optimal one in the underlying polyhedron (such as, e.g., relying on an extended formulation with one extra dimension to construct a pyramid whose apex is adjacent to all extreme points of the initial polyhedron). Such standard transformations usually make the polyhedron become highly degenerate. In presence of degeneracy, an extreme point might have exponentially many distinct bases defining it, and the simplex algorithm might perform an exponential number of consecutive degenerate pivots, staying in the same vertex, (see, e.g., [9]). Such behavior is referred to as stalling. In some pathological cases, certain pivot rules might even provide an infinite sequence of consecutive degenerate pivots at a degenerate vertex, a phenomenon called cycling. Although cycling can be easily avoided by employing, for instance, Bland’s rule [5] or a lexicographic rule [36], there is no known pivot rule that prevents stalling. As stated in several papers (e.g., [25, 26]), it is well known that solving a general LP in strongly polynomial time can be reduced to finding a pivot rule that prevents stalling in general polyhedra. However, there are a few cases in which it is known that the issue of stalling can be handled. The most famous example is the class of transportation polytopes, for which pivot rules with polynomial bounds on the number of consecutive degenerate pivots were introduced in [9] and further developed in [1, 16, 29]. See also the work of [28] for a strongly polynomial version of the primal network simplex.
Our findings are inspired by [18] to a great extent. The latter work shows that, given an LP in standard equality form with a totally unimodular coefficient matrix $A\in\mathbb{\{}-1,0,1\}^{m\times n}$, a vertex solution $x$ and an improving feasible circuit direction for $x$, one can construct a pivot rule which performs at most $m$ consecutive degenerate pivots. We generalize their proof, and show that one can employ any improving direction at a vertex $x$ of a general polyhedron, to avoid stalling. Our result, proved in Section 2, is summarized in the following theorem.
Theorem 1.
Given any LP of the form $\min\{c^{T}x:Ax=b,x\geq 0\}$ with $A\in\mathrm{\hskip 0.28453ptIR}^{m\times n}$, there exists a pivot rule that limits the number of consecutive degenerate simplex pivots at any non-optimal extreme solution to at most $n-m-1$.
We stress here that the pivots considered in Theorem 1 are degenerate simplex pivots, meaning that each of them yields an improving direction at the given extreme point, though this direction is not feasible. It is important to point this out as in general, given two adjacent extreme points $x,x^{\prime}$, one can easily perform a sequence of basis exchanges that yield $x^{\prime}$ from $x$ by identifying the common tight linearly independent constraints, and introducing each of them to the current basis in any order, until the direction $x^{\prime}-x$ is seen. However, we here want a strategy that guarantees that each basis exchange will define an improving direction.
We then discuss some byproducts of our result in Section 3. In particular, we revise the analysis of the simplex algorithm by Kitahara and Mizuno [19, 20] who bound the number of non-degenerate pivots it in terms of $n,m$ and the maximum and the minimum non-zero coordinate of a basic feasible solution. Their analysis combined with our degeneracy-escaping technique show that the total number of simplex pivots (both degenerate and non-degenerate) can be bounded in a similar way. As a consequence, one can have a strongly-polynomial number of simplex pivots for several combinatorial LPs (see observations following Corollary 6).
Of course the drawback of the whole machinery is that it requires an improving feasible direction at a given vertex. Though it is efficiently computable, this is in general as hard as solving an LP. For some classes of polytopes though (e.g., matching or flow polytopes) finding such a direction can be easier, thus making it worthwhile to apply our pivot rule. Most importantly, we think that the main importance of our result is from a theoretical perspective: it shows that, for several polytopes, not only a short path on the 1-skeleton exists, but a short sequence of simplex pivots always exists (and can also be efficiently computed). That is, a short sequence of basis exchanges that follow improving directions, when performing both degenerate and non-degenerate pivots.
2 Antistalling pivot rule
The goal of this section is to prove Theorem 1. Before doing that, we state some preliminaries, and also give a geometric intuition behind the result.
2.1 Preliminaries
The simplex algorithm works with LP in standard equality form:
$\begin{array}[]{rl}\min&c^{T}x\\ s.t.&Ax=b\\ &x\geq 0\end{array}$ (1)
For $A\in\mathrm{\hskip 0.28453ptIR}^{m\times n}$ and $B\subseteq[n]$ we use $A_{B}$ to denote the submatrix of $A$ formed by the columns indexed by $B$. Similarly, for $x\in\mathrm{\hskip 0.28453ptIR}^{n}$ and $B\in[n]$, the notation $x_{B}$ is used to denote a vector consisting of the entries of $x$ indexed by $B$. A basis of (1) is a subset $B\subseteq[n]$ with $|B|=m$ and $A_{B}$ being non-singular. The point $x$ with $x_{B}=A_{B}^{-1}b$, $x_{N}=\mathbbm{O}$ where $N:=[n]\setminus B$ is a basic solution of (1) with basis $B$. If additionally $x_{B}\geq 0$, both the basic solution $x$ and the basis $B$ are feasible. If $x_{i}>0$ for all $i\in B$, then $B$ and $x$ are called non-degenerate and degenerate otherwise, i.e., if $x_{i}=0$ for some $i\in B$. We let $\overline{A}:=A_{B}^{-1}A_{N}$ and $\overline{c}_{N}^{T}:=c_{N}^{T}-c_{B}^{T}\overline{A}$. In particular, $\overline{c}_{N}\in\mathrm{\hskip 0.28453ptIR}^{N}$ is the vector of so-called reduced costs for the basis $B$. The coordinates of the reduced cost vector will be addressed as $\overline{c}_{N,i}$, where the first subscript will be dropped if the basis $B=[n]\setminus N$ is clear from the context.
The simplex algorithm considers in each iteration a feasible $B$. If all elements of $\overline{c}_{N}$ are non-negative, then the basis $B$ and the corresponding basic feasible solution $x$ are known to be optimal. Otherwise, the algorithm pivots by choosing a non-basic coordinate with negative reduced cost to enter the basis, say $f$. It then performs a minimum ratio test to compute an index $i$ that minimizes $\frac{x_{i}}{\overline{A}_{if}}$ among all indices $i\in B$ for which $\overline{A}_{if}>0$. Such index $i$ will be the one leaving the basis, and it corresponds to a basic coordinate which hits its bound first when moving along the direction given by the tight constraints indexed by $B\setminus\{f\}$. At each iteration there could be multiple candidate indices for entering the basis (all the ones with negative reduced cost), as well as multiple candidate indices to leave the basis (all the ones for which the minimum ratio test value is attained). The choice of the entering and the leaving coordinates is the essence of a pivot rule (see [35] for a survey). A pivot is degenerate if the attained minimum ratio test value is 0 (hence, the extreme point solution $x$ does not change), and non-degenerate otherwise. Note that only coordinates with negative reduced cost are considered for entering the basis, since only such a choice guarantees the simplex algorithm to make progress when a non-degenerate pivot occurs. To later emphasize that we only refer to that kind of improving pivots, we will call them simplex pivots.
Finally, we let $\mathrm{kern}(A):=\{y\in\mathrm{\hskip 0.28453ptIR}^{n}:Ay=0\}$. Given (1), we call a vector $y\in\mathrm{kern}(A)$ with $c^{T}y<0$ an improving direction. Such $y$ is said to be feasible at a basic feasible solution $x$ if $x+\epsilon y\geq 0$ for sufficiently small positive $\epsilon$. Note that for any $y\in\mathrm{kern}(A)$ and any basis $B$ of $A$, the following holds:
$c^{T}y=c^{T}_{B}y_{B}+c^{T}_{N}y_{N}=c^{T}_{B}(-\overline{A}y_{N})+c^{T}_{N}y_{N}=\overline{c}_{N}^{T}y_{N}$ (2)
2.2 A geometric intuition
[…]
2.3 Proof of Theorem 1
Proof.
Let $x$ be a degenerate basic feasible solution of (1) with basis $B$. Without loss of generality, assume $B=[m]$. Then, $x_{i}=0$ for $i\in N=[n]\setminus[m]$. Assume also that $x_{i}=0$ for $i\in[k]$ with $1\leq k\leq m$ and $x_{i}>0$ for $i=k+1,\dots,m$. Let $S(B):=\{i\in N\mid\overline{c}_{i}<0\}$. Since $x$ is not optimal, $S(B)\neq\emptyset$.
It is well known, that a basic feasible solution $x$ of (1) with basis $B$ is not optimal if and only if there exists an improving feasible direction, i.e., there exists $y^{0}\in\mathrm{\hskip 0.28453ptIR}^{n}$ such that $Ay^{0}=0$, $c^{T}y^{0}<0$ and $y^{0}_{i}\geq 0$ for all $i\in[k]\cup\big{(}[n]\setminus[m]\big{)}$. Consider any such $y^{0}$. Let $Q_{1}(y^{0},B):=\{i\in[k]\mid y^{0}_{i}>0\}$ and $Q_{2}(y^{0},B):=\{i\in[n]\setminus[m]\mid y^{0}_{i}>0\}$. Without loss of generality, let $Q_{1}(y^{0},B)=[r]$ with $0\leq r\leq k$ and $Q_{2}(y^{0},B)=\{m+1,\dots,m+t\}$ with $1\leq t\leq n-m$. See Table 1 for an illustration.
1 $r$ $k$ $m$ $f$ $m+t$ $n$
$x$ 0 0 0 0 0 + + 0 0 0 0 0
$y^{0}$ + + + 0 0 $\star$ $\star$ + + + 0 0
$z$star $\star$ $0_{+}$ $0_{+}$star $\star$ 1 0 0 0 0
$y^{0}+\alpha z$ 0 $0_{+}$ $0_{+}$ $0_{+}$ $0_{+}$star $\star$ + + + 0 0
Table 1: An illustration of entries of $x,y^{0},z$ and $y^{1}:=y^{0}+\alpha z$. A positive, a non-negative, a negative, and a sign-arbitrary entry is denoted by $+$, $0_{+}$, -- and $\star$, respectively. Without loss of generality, we here assumed that the entering index $f$ is $m+1$, and that an index $g$ for which the minimum in Case III is attained is 1.
Since $c^{T}y^{0}=\overline{c}_{N}^{T}y_{N}^{0}<0$ due to (2), it follows that $S(B)\cap Q_{2}(y^{0},B)\neq\emptyset$. We choose the entering index to be a non-basic one with the most negative reduced cost [^3] We would like to point out that to prove the theorem, it suffices to choose any non-basic index with negative reduced cost in the support of $y^{0}$ as the entering index. The fact that $f$ has the most negative reduced cost among such indices will only be used for the results in the next section. in the support of $y^{0}$, that is $f=\mathrm{argmin}_{i\in S(B)\cap Q_{2}(y^{0},B)}\overline{c}_{i}$. To detect the leaving index, we consider the following case distinction. Note that the case distinction depends on the basis $B$ and the improving feasible direction $y^{0}$.
Case I: There is no $i\in[k]$ with $\overline{A}_{if}>0$. In this case, the minimum ratio test for $f$ is strictly positive, that is:
(*$(*)\quad\min_{i\in Eq^{>}(f)}\frac{x_{i}}{\overline{A}_{if}}>0\mbox{ where }Eq^{>}(f):=\{i\in[m]\mid\overline{A}_{if}>0\}\,.$
In particular, this means $Eq^{>}(f)\subseteq[m]\setminus[k]$. We perform a non-degenerate pivot, by selecting an index that minimizes (*)(*)( * ) as the leaving index.
Case II: $\overline{A}_{if}>0$ for some $i\in[k]\setminus[r]$. In this case, we perform a degenerate pivot by selecting $i$ as the leaving index. Let $B^{\prime}:=B\cup\{f\}\setminus\{i\}$. Because of degeneracy, the basic solution associated with $B^{\prime}$ is still $x$, and hence $y^{0}$ is still an improving feasible direction for $x$. Note that $|Q_{2}(y^{0},B^{\prime})|=|Q_{2}(y^{0},B)|-1$— since $y^{0}_{f}>0$ but $y^{0}_{i}=0$ by definition. Repeat the same for $B^{\prime}$ and $y^{0}$.
Case III: $\overline{A}_{if}\leq 0$ for all $i\in[k]\setminus[r]$ and $\overline{A}_{if}>0$ for some $i\in[r]$. In this case, we are going to change our improving feasible direction. Consider the following vector $z\in\mathrm{\hskip 0.28453ptIR}^{n}$ which is, in fact, an improving (thought not feasible) direction for the entering variable $x_{f}$:
$z_{i}=\begin{cases}-\overline{A}_{if},&\text{for $i\in[m]$}\\ 1,&\text{for $i=f$}\\ 0,&\text{otherwise}\end{cases}$ (3)
Note that $Az=0$ and $c^{T}z=\overline{c}_{f}<0$. Moreover $z_{i}\geq 0$ for each $i\in\big{(}[r]\setminus Eq^{>}(f)\big{)}\cup\big{(}[k]\setminus[r]\big{)}\cup\big{(}[n]\setminus[m]\big{)}$ and $z_{i}<0$ for $i\in Eq^{>}(f)$ by definition. Set
$y^{1}:=y^{0}+\alpha z\;\mbox{ with }\;\alpha:=\min_{i\in Eq^{>}(f)\cap[r]}\frac{y_{i}^{0}}{|z_{i}|}>0\,.$
Observe that $y^{1}$ is a feasible direction for $x$, since $A(y+\alpha z)=0$, and $y^{1}_{i}\geq 0$ for $i\in[k]\cup\big{(}[n]\setminus[m]\big{)}$, because of the choice of $z$ and $\alpha$. Furthermore, $y^{1}$ is an improving direction, because $c^{T}z<0$ and hence
$c^{T}y^{1}=c^{T}y^{0}+\alpha c^{T}z<c^{T}y^{0}<0\,.$ (4)
Note that $Q_{2}(y^{1},B)=Q_{2}(y^{0},B)$. Furthermore, let $g\in Eq^{>}(f)\cap[r]$ be the index for which the value of $\alpha$ is attained. We have $y^{1}_{g}=0$ and $\overline{A}_{gf}>0$. We now repeat the case distinction above for $B$ and $y^{1}$.
The key observation is that when the third case of the above case distinction has occurred, repeating the same for $B$ and $y^{1}$ falls into the second case (because the basis $B$ has not changed, meanwhile $y^{1}_{g}=x_{g}=0$ with $g\in B$ and $\overline{A}_{gf}>0$). Therefore, after each pivot, the cardinality of the support of the improving feasible direction in the non-basic indices decreases. Hence, a sequence of $|Q_{2}(y,B)|$ degenerate pivots would yield an improving feasible direction $y^{\prime}$ and a basis $B^{\prime}$ of $x$ with $Q_{2}(y^{\prime},B^{\prime})=0$, which in turn implies $0=\overline{c}^{T}_{N^{\prime}}y^{\prime}_{N^{\prime}}=c^{T}y^{\prime}$ due to (2), yielding a contradiction. Hence the number of consecutive degenerate pivots with the suggested pivot rule cannot exceed $n-m-1$.
3 Exploiting the antistalling rule for general bounds
Here we combine the result of the previous section with the analysis of [19, 20]. The authors of the latter works give a bound on the number of non-degenerate simplex pivots that depends on $n,m$ and the maximum and the minimum non-zero coordinate of basic solutions. The combined analysis yields a similar bound on the total number of (both degenerate and non-degenerate) simplex pivots.
We need a few additional notations. For any vector $x$, we let $\mathrm{supp}(x)$ denote its support. We denote the smallest and the largest non-zero coordinate of any basic feasible solution of (1) by $\delta$ and $\Delta$, respectively. We let $x^{\star}$ be an optimal basic feasible solution of (1). Finally, for a generic iteration $q$ of the simplex algorithm with basis $B^{q}$ and an improving feasible direction $y^{q}$, we let $\Delta^{q}_{\overline{c}}:=\max_{\{i\in N^{q}\mid y^{q}_{i}>0\}}-\overline{c}_{N^{q},i}$. Note that $\Delta^{q}_{\overline{c}}$ is the absolute value of the reduced cost of the entering variable according to the antistalling pivot rule defined in the previous section when using $y^{q}$.
We make use of the following result from [20].
Lemma 2 (Lemma 4 of [20]).
If these exists a constant $\lambda>0$ such that
$c^{T}(x^{q+1}-x^{\star})\leq\big{(}1-\frac{1}{\lambda}\big{)}c^{T}(x^{q}-x^{\star})$ (5)
holds for any consecutive distinct basic feasible solutions $x^{q}\neq x^{q+1}$ generated by the simplex algorithm (with any pivot rule), the total number of distinct basic feasible solutions encountered is at most
$(n-m)\Big{\lceil}\lambda\log_{e}\frac{m\Delta}{\delta}\Big{\rceil}\,.$
Given (1), apply the simplex algorithm with the antistalling pivot rule described in the previous section. In particular, at a general iteration $t$ of the algorithm, let $B^{t},x^{t}$, and $y^{t}$ be respectively the basis, the basic solution, and the improving feasible direction considered by the algorithm. Let $N^{t}:=[n]\setminus B^{t}$. If $x^{t}\neq x^{t-1}$ (i.e., we encounter $x^{t}$ for the first time, as a result of a non-degenerate pivot) let $y^{t}:=x^{\star}-x^{t}$. Observe that, once this is specified, the improving feasible direction is determined by the antistalling rule for all remaining degenerate pivots at $x^{t}$.
Lemma 3.
Let $x^{t}=x^{t+k}$ be the basic feasible solution associated with bases $B^{t}$ and $B^{t+k}$, and assume $x^{t+k}\neq x^{t+k+1}$. The following holds:
  • (a)
    $c^{T}y^{t+k}\leq c^{T}y^{t}$.
  • (b)
    $y^{t+k}_{i}=y^{t}_{i}$ for all $i\in N^{t+k}\cap\,\mathrm{supp}(y^{t+k})$.
Proof.
The statement in (a) follows from (4). The statement in (b) follows from the construction of our antistalling rule by induction: if Case III never occurs during the $k$ degenerate pivots, then $y^{t+k}=y^{t}$ and the statement holds trivially. Suppose that the situation of Case III appears for $B^{t+j},y^{t+j}$. The improving feasible direction changes (as $y^{t+j+1}=y^{t+j}+\alpha z$ ), but among the non-basic coordinates this only affects the value of $y^{t+j+1}_{f}$ (where $f$ is the entering index). Immediately after, Case II occurs and $f$ gets pivoted in, while a basic index $i$ with $y^{t+j+1}_{i}=0$ gets pivoted out, i.e., $B^{t+j+1}=\big{(}B^{t+j}\setminus\{i\}\big{)}\cup\{f\}$ and $N^{t+j+1}=\big{(}N^{t+j}\setminus\{f\}\big{)}\cup\{i\}$. Hence, the statement holds. ∎
The next result, which establishes (5) with $\lambda:=\frac{(n-m)\Delta}{\delta}$ for the simplex algorithm with the antistalling pivot rule, is inspired by [20][Theorem 3].
Lemma 4.
Let $x^{t}=x^{t+k}$ be the basic feasible solution associated with bases $B^{t}$ and $B^{t+k}$. Assume $x^{t-1}\neq x^{t}$ and $x^{t+k}\neq x^{t+k+1}$. We have
$c^{T}x^{t+k+1}-c^{T}x^{\star}\leq\big{(}1-\frac{\delta}{(n-m)\Delta}\big{)}(c^{T}x^{t+k}-c^{T}x^{\star})\,.$
Proof.
The optimality gap for $x^{t+k}$ can be bounded as follows:
$\begin{array}[]{rcl}c^{T}x^{t+k}-c^{T}x^{\star}&=&c^{T}x^{t}-c^{T}x^{\star}\\ &=&-c^{T}y^{t}\\ &\leq&-c^{T}y^{t+k}\\ &=&-\overline{c}^{T}_{N^{t+k}}y^{t+k}_{N^{t+k}}\\ &=&\sum_{i\in N^{t+k}}-\overline{c}^{T}_{N^{t+k},i}y^{t+k}_{i}\\ &=&\sum_{i\in N^{t+k}\mid y^{t+k}_{i}>0}-\overline{c}^{T}_{N^{t+k},i}y^{t+k}_{i}\\ &\leq&(n-m)\Delta^{t+k}_{\overline{c}}\Delta\,,\end{array}$ (6)
where we used Lemma 3(a) for the first inequality, and (2) for the third equality. The last equality follows from $y^{t+k}_{i}\geq 0,i\in N^{t+k}$ which is implied by feasibility of $y^{t+k}$ at $x^{t+k}$. The last inequality in turn follows from $\max_{\{i\in N^{t+k}\mid y^{t+k}_{i}>0\}}-\overline{c}_{N^{t+k},i}=\Delta^{t+k}_{\overline{c}}$, the fact that $|N^{t+k}|=n-m$, and since $y^{t+k}_{i}=y^{t}_{i}=x^{\star}_{i}-x^{t}_{i}\leq\Delta$ holds for any $i\in N^{t+k}\cap\,\mathrm{supp}(y^{t+k})$, using Lemma 3(b).
[…]
$\begin{array}[]{rl}c^{T}x^{t+k}-c^{T}x^{t+k+1}&=\overline{c}_{N^{t+k}}^{T}(x^{t+k}-x^{t+k+1})_{N^{t+k}}\\ &=-\overline{c}_{N^{t+k},f}x^{t+k+1}_{f}\\ &\geq\Delta^{t+k}_{\overline{c}}\delta\\ &\geq\frac{\delta}{(n-m)\Delta}c^{T}(x^{t+k}-x^{\star})\end{array}$
hold, where we used (2) for the first equality and (6) for the last inequality. Rearranging the terms yields the lemma statement. ∎
Now, combining Lemma 2 and Lemma 4 allows to conclude that the simplex algorithm with the antistalling pivot rule described in Section 2 encounters at most $(n-m)\lceil\frac{(n-m)\Delta}{\delta}\log_{e}\big{(}m\frac{\Delta}{\delta}\big{)}\rceil$ distinct basic feasible solutions. Since Theorem 1 in turn bounds the number of consecutive non-degenerate steps, combining it with the latter result yields a bound on the total number of pivots required to reach an optimal vertex. However, this bound does not take into account the number of (degenerate) pivots that might have to be performed at an optimal vertex to encounter a basis satisfying the optimality criterion. This is due to the fact that the antistalling pivot rule requires an improving feasible direction, which we do not have at an optimal vertex. Hence, we have to handle this case separately.
Let $B$ and $B^{\star}$ be a non-optimal and an optimal basis, both associated with an optimal basic feasible solution $x^{\star}$ of (1). Since $B$ is not optimal, there exists $f\in N$ with $\overline{c}_{N,f}=c^{T}z<0$ with $z$ as in (3). Observe that there exists $i\in B\cap N^{\star}$ with $\overline{A}_{if}>0$ and $x_{i}=0$: otherwise, all coordinates of $z_{N^{\star}}$ would be non-negative and hence $z\in C_{N^{\star}}:=\{x\in\mathrm{\hskip 0.28453ptIR}^{n}\mid Ax=0,x_{N^{\star}}\geq 0\}$. The latter however contradict the fact that all extreme rays of the above cone $C_{N^{\star}}$ have non-negative scalar products $\overline{c}_{N^{\star}}$ with $c$ due to optimality of $B^{\star}$. Hence one could perform a simplex pivot on $B$ with entering variable $f$ and leaving variable $i$. Let $B^{\prime}:=\big{(}B\setminus\{i\}\big{)}\cup\{f\}$ and $N^{\prime}:=\big{(}N\setminus\{f\}\big{)}\cup\{i\}$. Note that $\overline{c}_{N^{\prime},i}\geq 0$ and $i\in N^{\prime}\cap N^{\star}$. Either $B^{\prime}$ is a optimal basis and we stop, or there exists $j\in N^{\prime}\setminus\{i\}$ with $\overline{c}_{N^{\prime},j}<0$. In the latter case, however, we can enforce the constraint $x_{i}=0$ by removing the variable $x_{i}$ together with the corresponding column of $A$ and entry $c_{i}$ of $c$ from (1). By doing so we obtain a new LP with the number of variables smaller by one that has $B^{\prime}$ and $B^{\star}$ as a non-optimal and an optimal basis, respectively, since $\overline{c}_{N^{\prime}\setminus\{i\},j}<0$ and $\overline{c}_{N^{\star}\setminus\{i\}}\geq 0$. We can set $B=B^{\prime}$ and repeat the process. Since at each iteration a variable with index from $N^{\star}$ gets removed, after at most $n-m$ iterations $B=B^{\star}$. Since the number of degenerate pivots at an optimal vertex is bounded by $n-m$, we can state the following result.
Theorem 5.
Given an LP (1) and an initial feasible basis, there exists a pivot rule that makes the simplex algorithm reach an optimal basis in at most $(n-m)^{2}\lceil\frac{(n-m)\Delta}{\delta}\log_{e}\big{(}m\frac{\Delta}{\delta}\big{)}\rceil$ simplex pivots.
In general, the value of $\delta$ is NP-hard to compute [22]. However, observe that for integral $A$ and $b$ (which can be assumed without loss of generality for rational LPs), one can bound $\Delta\leq||b||_{1}\Delta_{A}$ and $\delta\geq\frac{1}{\Delta_{A}}$ due to Cramer’s rule, where $\Delta_{A}$ is the largest absolute value of the determinant of a square $m\times m$ submatrix of $A$. Then the next statement a straightforward corollary of the above theorem.
Corollary 6.
For any basic feasible solution of an LP (1) with integral $A$ and $b$, there exists a sequence of at most $(n-m)^{2}\lceil(n-m)\Delta_{A}^{2}||b||_{1}\log_{e}\big{(}m\Delta_{A}^{2}||b||_{1}\big{)}\rceil$ simplex pivots leading to an optimal basis.
We conclude this section by observing that, using the above corollary, one can prove existence of short sequences of simplex pivots (that is, of length strongly-polynomial in $n,m$) between any two extreme points of several combinatorial LPs. Just to mention a few:
(a) LPs modeling matching/vertex-cover/edge-cover/stable-set problems in bipartite graphs. For matchings, the LP maximizes a linear function over a set of constraint of the form $\{A^{\prime}x\leq 1,x\geq 0\}$, where the coefficient matrix $A^{\prime}$ is the node-edge incidence matrix of an undirected bipartite graph. After putting the LP in standard equality form by adding slack variables, we get constraints of the form $\{Ax=1,x\geq 0\}$, where $A$ is a totally unimodular matrix (and so $\Delta_{A}=1$). The result then follows from Corollary 6. The same holds for vertex-cover (constraints of the form $\{{A^{\prime}}^{T}x\geq 1,x\geq 0\}$), edge-cover (constraints of the form $\{{A^{\prime}}x\geq 1,x\geq 0\}$), stable-set (constraints of the form $\{{A^{\prime}}^{T}x\leq 1,x\geq 0\}$).
(b) LP relaxations for the above problems for general graphs (i.e., optimization over the fractional matching/vertex-cover/edge-cover/stable-set polytopes). The same LPs as in (a), in general graphs, have a constraint matrix $A^{\prime}$ (resp. ${A^{\prime}}^{T}$) that is not totally unimodular. However, the set of constraints defines half-integral polytopes (see [2, 27]). Note that, after putting the LPs in standard equality form, the slack variables can be loosely bounded from above by the number of variables $n$. Hence, $\Delta\leq n$ and $\delta\geq\frac{1}{2}$, and the result follows from Theorem 5.
(c) LP for the stable marriage problem [30, 38]. The (exact) LP formulation has constraints of the form $\{A^{\prime}x\leq 1,B^{\prime}x\geq 1,x\geq 0\}$. After putting the LP in standard equality form, the slack variables can be bounded by $1$. Hence, $\Delta=\delta=1$, and the result follows from Theorem 5.
(d) LPs modeling various flow problems (such as max flows, min cost flows, flow circulations) with unit (or bounded) capacity values. For the corresponding LPs (see [32]), the constraint matrix will be a node-arc incidence matrix of a directed graph, which is totally unimodular, and the right-hand-side vector will be bounded by the capacity values. Similarly to (a), one can rely on Corollary 6 to get the result.
Acknowledgements.
The authors are grateful to Volker Kaibel for helpful discussions on the geometric interpretation of the antistalling pivot rule. We thank Alex Black and anonymous reviewers for their helpful comments. The authors express gratitude to the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) for supporting the first author within 314838170, GRK 2297 MathCoRe and providing travel funds for visiting the second author in Milan. The second author is supported by the NWO VIDI grant VI.Vidi.193.087.