### v4

Abstract
The simplex algorithm is one of the most popular algorithms to solve linear programs (LPs). Starting at an extreme point solution of an LP, it performs a sequence of basis exchanges (called pivots) that allows one to move to a better extreme point along an improving edge-direction of the underlying polyhedron.
A key issue in the simplex algorithm’s performance is degeneracy, which may lead to a (potentially long) sequence of basis exchanges which do not change the current extreme point solution.
In this paper, we prove that one can employ any improving feasible direction at an extreme point to limit the number of consecutive degenerate pivots that the simplex algorithm performs to nm1n-m-1, where $n$ is the number of variables and $m$ is the number of equality constraints of a given LP in standard equality form.
1 Introduction
The simplex algorithm first introduced in [14] has always been one of the most widely used algorithms for solving linear programming problems (LPs). Although it performs as a linear time algorithm for many real world linear programming models (see, e.g., the survey [39]) and a variant of it (the shadow simplex) has been shown to have polynomial smoothed complexity (see, e.g., [7, 8, 13, 29, 40, 46]), no polynomial version of the simplex algorithm has been found until these days, despite decades of research.
The simplex algorithm relies on the concept of basis, where a basis corresponds to an inclusion-wise minimal set of tight constraints which defines an extreme point solution of the underlying linear program. In each step, it performs a basis exchange by replacing one constraint currently in the basis with a different one that is not in the basis. Such an exchange is called a pivot. From a geometric perspective, a basis exchange identifies a direction to move from the current extreme point. The simplex algorithm only considers pivots that yield an improving direction, with respect to the objective function to be optimized. As the result of pivoting, it is possible for the algorithm to either stay at the same extreme point, or to move to an adjacent extreme point of the underlying polyhedron. We refer to a pivot of the former type as degenerate, in contrast to a pivot of the latter type, which will be called non-degenerate. The pivot rule which determines how to perform these basis exchanges (i.e., deciding which inequality enters and which inequality leaves the basis) is the core of the simplex algorithm. Various pivot rules have been introduced over decades, but none of them yields a polynomial time version of the simplex algorithm so far (see, e.g., [3, 16, 20, 18, 21, 41, 26, 47] and the references therein).
When analyzing the performance of the simplex algorithm, one can often assume that the underlying polyhedron $\mathcal{P}$ is non degenerate, e.g., by performing a small perturbation of the constraint’s right-hand-side. In this case, a necessary condition to guarantee an efficient performance of the simplex algorithm, is that there is always a way to reach an optimal solution within a polynomial number of non-degenerate pivots. Geometrically, this would imply that there is always a path between two extreme points on the 1-skeleton of $\mathcal{P}$ (the graph where vertices correspond to the extreme points of $\mathcal{P}$, and edges correspond to the 1-dimensional faces of $\mathcal{P}$), whose length is bounded by a polynomial in the input description of $\mathcal{P}$. This is a major and still open question in the field, commonly referred to as the polynomial Hirsch Conjecture (see for example [6, 12, 27, 37, 43]).
On the other hand, even in the case when a short path exists, it is not clear how the simplex algorithm would be bound to follow it (see for instance the case of $0/1$-polytopes discussed in [4], or the hardness results in [15, 9]). In fact, given any LP one can apply some transformation to it to ensure the existence of a short path between an initial extreme point solution and an optimal one in the underlying polyhedron (such as, e.g., relying on an extended formulation with one extra dimension to construct a pyramid whose apex is adjacent to all extreme points of the initial polyhedron). Such standard transformations usually make the polyhedron become highly degenerate. In the presence of degeneracy, an extreme point might have exponentially many distinct bases defining it, and the simplex algorithm might perform an exponential number of consecutive degenerate pivots, staying in the same vertex, (see, e.g., [11]). Such behavior is referred to as stalling. In some pathological cases, certain pivot rules might even provide an infinite sequence of consecutive degenerate pivots at a degenerate vertex, a phenomenon called cycling. Although cycling can be easily avoided by employing, for instance, Bland’s rule [5] or a lexicographic rule [42], there is no known pivot rule that prevents stalling. As stated in several papers (e.g., [28, 30]), it is well known that solving a general LP in strongly polynomial time can be reduced to finding a pivot rule that prevents stalling in general polyhedra, and that can be implemented in strongly polynomial time. However, there are a few cases in which it is known that the issue of stalling can be handled. The most famous example is the class of transportation polytopes, for which pivot rules with polynomial bounds on the number of consecutive degenerate pivots were introduced in [11] and further developed in [1, 19, 35]. See also the work of [33] for a strongly polynomial version of the primal network simplex.
Our findings are inspired by [22] to a great extent. The latter work shows that, given an LP in standard equality form with a totally unimodular coefficient matrix $A\in\mathbb{\{}-1,0,1\}^{m\times n}$, a vertex solution $x$ and an improving feasible circuit direction for $x$, one can construct a pivot rule which performs at most $m$ consecutive degenerate pivots. We generalize their proof, and show that one can employ any improving feasible direction at a vertex $x$ of a general polyhedron, to avoid stalling. Our result, proved in Section 2, is summarized in the following theorem.
Theorem 1.1
Given any LP of the form $\min\{c^{T}x:Ax=b,x\geq 0\}$ with $A\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{m\times n}$, there exists a pivot rule that limits the number of consecutive degenerate simplex pivots at any non-optimal extreme solution to at most nm1n-m-1.
The pivot rule can be efficiently implemented whenever an improving feasible direction at a non-optimal extreme point is available.
We stress here that the pivots considered in Theorem 1.1 are degenerate simplex pivots, meaning that each of them yields an improving direction at the given extreme point, though this direction is not feasible. It is important to point this out as in general, given two adjacent extreme points $x,x^{\prime}$, one can easily perform a sequence of basis exchanges that yields $x^{\prime}$ from $x$ by identifying the common tight linearly independent constraints, and introducing each of them to the current basis in any order, until the direction $x^{\prime}-x$ is seen. However, we here want a strategy that guarantees that each basis exchange defines an improving direction, and hence is realizable by the simplex algorithm.
We then discuss some byproducts of our result in Section 3. In particular, we revise the analysis of the simplex algorithm by Kitahara and Mizuno [24, 25] who bound the number of non-degenerate pivots in terms of $n,m$ and the maximum and the minimum non-zero coordinate of a basic feasible solution. Their analysis combined with our degeneracy-escaping technique show that the total number of simplex pivots (both degenerate and non-degenerate) can be bounded in a similar way. As a consequence, one can have a strongly-polynomial number of simplex pivots for several combinatorial LPs. In addition, we perform some computational experiments to evaluate the performance of the antistalling pivot rule in practice, reported in Section 4.
Of course the drawback of the whole machinery is that it requires an improving feasible direction at a given vertex. Though it is efficiently computable, this is in general as hard as solving an LP. For some classes of polytopes though (e.g., matching or flow polytopes) finding such a direction can be easier, thus making it worthwhile to apply our pivot rule. Most importantly, we think that the main importance of our result is from a theoretical perspective: it shows that, for several polytopes, not only a short path on the 1-skeleton exists, but a short sequence of simplex pivots always exists (and can also be efficiently computed). That is, a short sequence of basis exchanges that follow improving directions, when performing both degenerate and non-degenerate pivots.
2 Antistalling pivot rule
The goal of this section is to prove Theorem 1.1. Before doing that, we state some preliminaries, and also give a geometric intuition behind the result.
2.1 Preliminaries
The simplex algorithm works with LPs in standard equality form:
$\begin{array}[]{rll}\min&c^{T}x&\\ \mbox{s.t.}&Ax&=b\\ &x&\geq 0\end{array}$ (1)
For $A\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{m\times n}$ and $B\subseteq[n]{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}:=\{1,2,\dots,n\}}$, we use $A_{B}$ to denote the submatrix of $A$ formed by the columns indexed by $B$. Similarly, for $x\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{n}$ and $B\in[n]$, the notation $x_{B}$ is used to denote a vector consisting of the entries of $x$ indexed by $B$. A basis of (1) is a subset $B\subseteq[n]$ with $|B|=m$ and $A_{B}$ being non-singular. The point $x$ with $x_{B}=A_{B}^{-1}b$, $x_{N}=0$ where $N:=[n]\setminus B$ is a basic solution of (1) with basis $B$. If additionally $x_{B}\geq 0$, both the basic solution $x$ and the basis $B$ are feasible. If $x_{i}>0$ for all $i\in B$, then $B$ and $x$ are called non-degenerate. If instead $x_{i}=0$ for some $i\in B$, then $B$ and $x$ are called degenerate. We let $\overline{A}:=A_{B}^{-1}A_{N}$ and $\overline{c}_{N}^{T}:=c_{N}^{T}-c_{B}^{T}\overline{A}$. In particular, $\overline{c}_{N}\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{N}$ is the vector of so-called reduced costs for the basis $B$. The coordinates of the reduced cost vector will be addressed as $\overline{c}_{N,i}$, where the first subscript will be dropped if the basis $B=[n]\setminus N$ is clear from the context.
The simplex algorithm considers in each iteration a feasible $B$. If all elements of $\overline{c}_{N}$ are non-negative, then the basis $B$ and the corresponding basic feasible solution $x$ are known to be optimal. Otherwise, the algorithm pivots by choosing a non-basic coordinate with negative reduced cost to enter the basis, say $f$. It then performs a minimum ratio test to compute an index $i$ that minimizes $\frac{x_{i}}{\overline{A}_{if}}$ among all indices $i\in B$ for which $\overline{A}_{if}>0$. Such index $i$ will be the one leaving the basis, and it corresponds to a basic coordinate which hits its non-negativity bound first when moving along the direction given by the tight constraints indexed by $B\setminus\{f\}$. At each iteration there could be multiple candidate indices for entering the basis (all the ones with negative reduced cost), as well as multiple candidate indices to leave the basis (all the ones for which the minimum ratio test value is attained). The choice of the entering and the leaving coordinates is the essence of a pivot rule (see [41] for a survey). A pivot is degenerate if the attained minimum ratio test value is 0 (hence, the extreme point solution $x$ does not change), and non-degenerate otherwise. Note that only coordinates with negative reduced cost are considered for entering the basis, since only such a choice guarantees the simplex algorithm to make progress when a non-degenerate pivot occurs. To later emphasize that we only refer to that kind of improving pivots, we will call them simplex pivots.
Finally, we let $\mathrm{kern}(A):=\{y\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{n}:Ay=0\}$. Given (1), we call a vector $y\in\mathrm{kern}(A)$ with $c^{T}y<0$ an improving direction. Such $y$ is said to be feasible at a basic feasible solution $x$ if $x+\varepsilon y\geq 0$ for sufficiently small positive $\varepsilon$. Note that for any $y\in\mathrm{kern}(A)$ and any basis $B$ of $A$, the following holds:
$c^{T}y=c^{T}_{B}y_{B}+c^{T}_{N}y_{N}=c^{T}_{B}(-\overline{A}y_{N})+c^{T}_{N}y_{N}=\overline{c}_{N}^{T}y_{N}$ (2)
2.2 A geometric intuition
Before providing a formal proof, we give a geometric intuition on how our pivot rule works. For this, it will be easier to abandon the standard equality form. In particular, consider a degenerate vertex $x\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{d}$ of a polytope $\mathcal{P}$ of dimension $d\in\mathbbm{N}$ defined by inequalities $a_{i}^{T}x\leq b_{i}$ with $a_{i}\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{d},b_{i}\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR},i\in[n]$. In this setting, degeneracy means that more than $d$ inequalities are tight at $x$. Consider a subset $N$ of the set of indices of all inequalities that are tight at $x$ with $|N|=d$ and let $B:=[n]\setminus N$ (see Figure 1). Note that here we intentionally redefine a notation used in the standard equality form, where non basic coordinates always have their corresponding constraints tight. Since $x$ is degenerate, there is at least one inequality that is tight at $x$ with index in $B$. We denote the subset of such inequalities by $W\subseteq B$. Finally, assume $x$ is not an optimal vertex of $\mathcal{P}$ when minimizing an objective function $c\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{d}$ over $\mathcal{P}$, and let $y^{0}\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{d}$ be an improving feasible direction at $x$, i.e., such that $x+\varepsilon y^{0}\in\mathcal{P}$ for a sufficiently small positive $\varepsilon$ and $c^{T}y^{0}<0$. In order to find an improving edge of $\mathcal{P}$ at $x$, we look at the directions given by the extremal rays of the basic cone $C(N):=\{x\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{n}\mid a_{i}^{T}x\leq 0,i\in N\}$. If there is an improving feasible direction among them, we are done. Otherwise, we reduce the dimension of the polytope in the following way. We pick an extremal ray $z\in C(N)$ such that $c^{T}z<0$. Note that $z$ is formed by d1d-1 inequalities from $N$. Let $f\in N$ be the only inequality index not used to define $z$, among the $d$ tight inequalities in $C(N)$ ($f$ is depicted as the gray fading facet in Figure 1). Consider a vector combination $y^{0}+\alpha z$ where $\alpha\geq 0$ and note that it provides an improving direction for any positive $\alpha$. Since $y^{0}$ is contained in the feasible cone $C(N\cup W)=\{x\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{n}\mid a_{i}^{T}x\leq 0,i\in N\cup W\}$ but $z$ is not, the vector $y^{0}+\alpha z$ leaves $C(N\cup W)$ for sufficiently large $\alpha$. Hence there has to be a number $\alpha^{1}\geq 0$ such that $y^{0}+\alpha^{1}z$ is contained in a facet of $C(N\cup W)$. Without loss of generality assume that the latter facet is defined by the $g\textsuperscript{th}$ inequality (green facet in Figure 1). Note that $g\in W$ since $y^{0}$ and $z$ are both contained in the basic cone $C(N)$ and so is $y^{0}+\alpha z$ for any non-negative $\alpha$. Then, we define a new improving feasible direction $y^{1}:=y^{0}+\alpha^{1}z$, $N=\big(N\setminus\{f\}\big)\cup\{g\}$, $B=\big(B\setminus\{g\}\big)\cup\{f\}$ and continue searching for an improving edge inside the facet defined by the $g\textsuperscript{th}$ inequality. Since $x+\varepsilon y^{1}$ belongs to this facet for an $\varepsilon>0$ small enough, the latter yields dimension reduction. After at most d1d-1 such steps we are bound to encounter a one-dimensional face of $\mathcal{P}$ containing $x+\varepsilon y^{\prime}$, where $y^{\prime}$ is an improving feasible direction at $x$. For an illustration see Figure 1.
Figure 1: To the left, facets corresponding to inequalities in NN and $W$ are colored in navy and white, respectively. To the right, the gray fading facet is defined by the $f\textsuperscript{th}$ inequality and the green one corresponds to the $g\textsuperscript{th}$.
2.3 Proof of Theorem 1.1
Proof
Let $x$ be a degenerate basic feasible solution of (1) with basis $B$. Without loss of generality, assume $B=[m]$. Then, $x_{i}=0$ for $i\in N=[n]\setminus[m]$. Assume also that $x_{i}=0$ for $i\in[k]$ with $1\leq k\leq m$ and $x_{i}>0$ for $i=k+1,\dots,m$. Let $S(B):=\{i\in N\mid\overline{c}_{i}<0\}$. Since $x$ is not optimal, $S(B)\neq\emptyset$.
It is well known that a basic feasible solution $x$ of (1) with basis $B$ is not optimal if and only if there exists an improving feasible direction, i.e., there exists $y^{0}\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{n}$ such that $Ay^{0}=0$, $c^{T}y^{0}<0$ and $y^{0}_{i}\geq 0$ for all $i\in[k]\cup\big([n]\setminus[m]\big)$. Consider any such $y^{0}$. Let $Q_{1}(y^{0},B):=\{i\in[k]\mid y^{0}_{i}>0\}$ and $Q_{2}(y^{0},B):=\{i\in[n]\setminus[m]\mid y^{0}_{i}>0\}$. Without loss of generality, let $Q_{1}(y^{0},B)=[r]$ with $0\leq r\leq k$ and $Q_{2}(y^{0},B)=\{m+1,\dots,m+t\}$ with $1\leq t\leq n-m$. See Table 1 for an illustration.
1 $r$ $k$ $m$ $f$ $m+t$ $n$
$x$ 0 0 0 0 0 + + 0 0 0 0 0
$y^{0}$ + + + 0 0 $\star$ $\star$ + + + 0 0
$z$ $\star$ $\star$ $_{+} 0+0_{+}$ $\star$ $\star$ 1 0 0 0 0
$y^{0}+\alpha z$ 0 $_{+} 0+0_{+}$ $_{+} 0+0_{+}$ $\star$ $\star$ + + + 0 0
Table 1: An illustration of entries of $x,y^{0},z$ and $y^{1}:=y^{0}+\alpha z$. A positive, a non-negative, a negative, and a sign-arbitrary entry is denoted by ++, $0_{+}$, - and $\star$, respectively. Without loss of generality, we here assumed that the entering index $f$ is $m+1$, and that an index $g$ for which the minimum in Case III is attained is 1.
Since $c^{T}y^{0}=\overline{c}_{N}^{T}y_{N}^{0}<0$ due to (2), it follows that $S(B)\cap Q_{2}(y^{0},B)\neq\emptyset$. We choose the entering index to be a non-basic one with the most negative reduced cost [^1] We would like to point out that to prove the theorem, it suffices to choose any non-basic index with negative reduced cost in the support of $y^{0}$ as the entering index. The fact that $f$ has the most negative reduced cost among such indices will only be used for the results in the next section. in the support of $y^{0}$, that is $f=\mathrm{argmin}_{i\in S(B)\cap Q_{2}(y^{0},B)}\overline{c}_{i}$. To detect the leaving index, we consider the following case distinction. Note that the case distinction depends on the basis $B$ and the improving feasible direction $y^{0}$.
Case I: There is no $i\in[k]$ with $\overline{A}_{if}>0$. In this case, the minimum ratio test for $f$ is strictly positive, that is:
$(*)\quad\min_{i\in Eq^{>}(f)}\frac{x_{i}}{\overline{A}_{if}}>0\mbox{ where }Eq^{>}(f):=\{i\in[m]\mid\overline{A}_{if}>0\}\,.$
In particular, this means $Eq^{>}(f)\subseteq[m]\setminus[k]$. We perform a non-degenerate pivot, by selecting an index that minimizes ()(*) as the leaving index.
Case II: $\overline{A}_{if}>0$ for some $i\in[k]\setminus[r]$. In this case, we perform a degenerate pivot by selecting $i$ as the leaving index. Let $B^{\prime}:=B\cup\{f\}\setminus\{i\}$. Because of degeneracy, the basic solution associated with $B^{\prime}$ is still $x$, and hence $y^{0}$ is still an improving feasible direction for $x$. Note that $|Q_{2}(y^{0},B^{\prime})|=|Q_{2}(y^{0},B)|-1$ since $y^{0}_{f}>0$ but $y^{0}_{i}=0$ by definition. Repeat the same for $B^{\prime}$ and $y^{0}$.
Case III: $\overline{A}_{if}\leq 0$ for all $i\in[k]\setminus[r]$ and $\overline{A}_{if}>0$ for some $i\in[r]$. In this case, we are going to change our improving feasible direction. Consider the following vector $z\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{n}$ which is, in fact, an improving (though not feasible) direction for the entering variable $x_{f}$:
$z_{i}=\begin{cases}-\overline{A}_{if}&\text{for $i\in[m]$,}\\ 1&\text{for $i=f$,}\\ 0&\text{otherwise.}\end{cases}$ (3)
Note that $Az=0$ and $c^{T}z=\overline{c}_{f}<0$. Moreover $z_{i}\geq 0$ for each $i\in\big([r]\setminus Eq^{>}(f)\big)\cup\big([k]\setminus[r]\big)\cup\big([n]\setminus[m]\big)$ and $z_{i}<0$ for $i\in Eq^{>}(f)$ by definition. Set
$y^{1}:=y^{0}+\alpha z\;\mbox{ with }\;\alpha:=\min_{i\in Eq^{>}(f)\cap[r]}\frac{y_{i}^{0}}{-z_{i}}>0\,.$
Observe that $y^{1}$ is a feasible direction for $x$, since $A(y+\alpha z)=0$, and $y^{1}_{i}\geq 0$ for $i\in[k]\cup\big([n]\setminus[m]\big)$, because of the choice of $z$ and $\alpha$. Furthermore, $y^{1}$ is an improving direction, because $c^{T}z<0$ and hence
$c^{T}y^{1}=c^{T}y^{0}+\alpha c^{T}z<c^{T}y^{0}<0\,.$ (4)
Note that $Q_{2}(y^{1},B)=Q_{2}(y^{0},B)$. Furthermore, let $g\in Eq^{>}(f)\cap[r]$ be the index for which the value of $\alpha$ is attained. We have $y^{1}_{g}=0$ and $\overline{A}_{gf}>0$. We now repeat the case distinction above for $B$ and $y^{1}$.
The key observation is that when the third case of the above case distinction has occurred, repeating the same for $B$ and $y^{1}$ falls into the second case (because the basis $B$ has not changed, meanwhile $y^{1}_{g}=x_{g}=0$ with $g\in B$ and $\overline{A}_{gf}>0$). Therefore, after each degenerate pivot, the cardinality of the support of the improving feasible direction in the non-basic indices decreases. Hence, a sequence of $|Q_{2}(y^{0},B)|$ degenerate pivots would yield an improving feasible direction $y^{\prime}$ and a basis $B^{\prime}$ of $x$ with $Q_{2}(y^{\prime},B^{\prime})=0$, which in turn implies $0=\overline{c}^{T}_{N^{\prime}}y^{\prime}_{N^{\prime}}=c^{T}y^{\prime}$ due to (2), yielding a contradiction. Hence the number of consecutive degenerate pivots with the suggested pivot rule cannot exceed nm1n-m-1. ∎
For the sake of clarity, the antistalling pivot rule is formally summarized in Algorithm 1.
Algorithm 1 Antistalling pivot rule.
0: A basic feasible solution $x$ of (1) with basis $B=\{1,\dots,m\}$ and an improving feasible direction $y$.
1: Compute the matrix $\overline{A}:=A_{B}^{-1}A_{N}$ and the reduced cost vector $\overline{c}_{N}^{T}:=c_{N}^{T}-c_{B}^{T}\overline{A}$, where $N:=[n]\setminus B$.
2: Compute the entering index $f$ as $\mathrm{argmin}_{\{i\in N:\overline{c}_{i}<0\}\cap\mathrm{supp}(y)}\overline{c}_{i}$.
3:if $\overline{A}_{if}\leq 0$ then
4:  The optimal cost is unbounded: the algorithm terminates.
5:else
6:  Perform the minimum ratio test:
$r:=\min_{i\in\mathrm{supp}\big(\overline{A}_{if}\big)}\frac{x_{i}}{\overline{A}_{if}}$ and $J:=\mathrm{argmin}_{i\in\mathrm{supp}\big(\overline{A}_{if}\big)}\frac{x_{i}}{\overline{A}_{if}}$.
7:  if $r>0$ (case I) then
8:   Choose $j\in J$ as the leaving variable. Perform non-degenerate pivot: $B=B\cup\{f\}\setminus\{j\}$.
9:  else if $J\cap B\setminus\mathrm{supp}(y)\neq\emptyset$ (case II) then
10:   Choose $j\in J\cap B\setminus\mathrm{supp}(y)$ as the leaving variable. Perform degenerate pivot: $B=B\cup\{f\}\setminus\{j\}$. Go back to 1.
11:  else
12:   (case III) Compute improving (though not feasible) direction $z$ as:
$z_{i}=\begin{cases}-\overline{A}_{if}&\text{for $i\in B$,}\\ 1&\text{for $i=f$,}\\ 0&\text{otherwise.}\end{cases}$
Compute the new improving feasible direction $y:=y+\alpha z$, where $\alpha:=\min_{i\in\mathrm{supp}(\overline{A}_{if})\cap\mathrm{supp}(y)\setminus\mathrm{supp}(x)}\frac{y_{i}}{\overline{A}_{if}}$. Go back to 3.
13:  end if
14:end if
We conclude with a strengthening of the bound of Theorem 1.1 which will be useful in the next section.
Theorem 2.1
Given any LP of the form $\min\{c^{T}x:Ax=b,x\geq 0\}$ with $A\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{m\times n}$, there exists a pivot rule that limits the number of consecutive degenerate simplex pivots at any non-optimal extreme solution $x$ to at most $\min\{n-m-1,m-1\}$.
The pivot rule can be efficiently implemented whenever an improving feasible direction of the form $\tilde{x}-x$, for some extreme point solution $\tilde{x}$, is available.
Proof
Let $x$ be the currently considered basic feasible solution with basis $B$. If $y^{0}$ is chosen to be equal to $\tilde{x}-x$ for some basic feasible solution $\tilde{x}$ of our LP with $c^{T}\tilde{x}<c^{T}x$, then one can observe that $|Q_{2}(y^{0},B)|\leq m$ because $\tilde{x}$ has at most $m$ non-zero coordinates. By the proof of Theorem 1.1 above, the latter value strictly decreases with each degenerate step produced by the antistalling pivot rule. Hence, the total number of consecutive degenerate pivots at the vertex $x$ can be strengthened to be at most $\min\{n-m-1,m-1\}$. ∎
3 Exploiting the antistalling rule for general bounds
Here we combine the result of the previous section with the analysis of [24, 25]. The authors of the latter works give a bound on the number of non-degenerate simplex pivots that depends on $n,m$ and the maximum and the minimum non-zero coordinate of basic solutions. The combined analysis yields a similar bound on the total number of (both degenerate and non-degenerate) simplex pivots.
We need a few additional notations. For any vector $x$, we let $\mathrm{supp}(x)$ denote its support. We denote the smallest and the largest non-zero coordinate of any basic feasible solution of (1) by $\delta$ and $\Delta$, respectively. We let $x^{\star}$ be an optimal basic feasible solution of (1). Finally, for a generic iteration $q$ of the simplex algorithm with basis $B^{q}$ and an improving feasible direction $y^{q}$, we let $\Delta^{q}_{\overline{c}}:=\max_{\{i\in N^{q}\mid y^{q}_{i}>0\}}-\overline{c}_{N^{q},i}$. Note that $\Delta^{q}_{\overline{c}}$ is the absolute value of the reduced cost of the entering variable according to the antistalling pivot rule defined in the previous section when using $y^{q}$.
We make use of the following result from [25].
Lemma 1(Lemma 4 of [25])
If there exists a constant $\lambda>0$ such that
$c^{T}(x^{q+1}-x^{\star})\leq\big(1-\frac{1}{\lambda}\big)c^{T}(x^{q}-x^{\star})$ (5)
holds for any consecutive distinct basic feasible solutions $x^{q}\neq x^{q+1}$ generated by the simplex algorithm (with any pivot rule), the total number of distinct basic feasible solutions encountered is at most
$(n-m)\Big\lceil\lambda\ln\frac{m\Delta}{\delta}\Big\rceil\,.$
Given (1), apply the simplex algorithm with the antistalling pivot rule described in the previous section. In particular, at a general iteration $t$ of the algorithm, let $B^{t},x^{t}$, and $y^{t}$ be the basis, the basic solution, and the improving feasible direction considered by the algorithm, respectively. Let $N^{t}:=[n]\setminus B^{t}$. If $x^{t}\neq x^{t-1}$ (i.e., we encounter $x^{t}$ for the first time, as a result of a non-degenerate pivot) let $y^{t}:=x^{\star}-x^{t}$. Observe that, once this is specified, the improving feasible direction is determined by the antistalling rule for all remaining degenerate pivots at $x^{t}$.
Lemma 2
Let $x^{t}=x^{t+k}$ be the basic feasible solution associated with bases $B^{t}$ and $B^{t+k}$, and assume $x^{t+k}\neq x^{t+k+1}$. The following holds:
  • (a)
    $c^{T}y^{t+k}\leq c^{T}y^{t}$.
  • (b)
    $y^{t+k}_{i}=y^{t}_{i}$ for all $i\in N^{t+k}\cap\,\mathrm{supp}(y^{t+k})$.
Proof
The statement in (a) follows from (4). The statement in (b) follows from the construction of our antistalling rule by induction: if Case III never occurs during the $k$ degenerate pivots, then $y^{t+k}=y^{t}$ and the statement holds trivially. Suppose that the situation of Case III appears for $B^{t+j},y^{t+j}$. The improving feasible direction changes (as $y^{t+j+1}=y^{t+j}+\alpha z$ ), but among the non-basic coordinates this only affects the value of $y^{t+j+1}_{f}$ (where $f$ is the entering index). Immediately after, Case II occurs and $f$ gets pivoted in, while a basic index $i$ with $y^{t+j+1}_{i}=0$ gets pivoted out, i.e., $B^{t+j+1}=\big(B^{t+j}\setminus\{i\}\big)\cup\{f\}$ and $N^{t+j+1}=\big(N^{t+j}\setminus\{f\}\big)\cup\{i\}$. Hence, the statement holds. ∎
The next result, which establishes (5) with $\lambda:=\frac{(n-m)\Delta}{\delta}$ for the simplex algorithm with the antistalling pivot rule, is inspired by [25][Theorem 3].
Lemma 3
Let $x^{t}=x^{t+k}$ be the basic feasible solution associated with bases $B^{t}$ and $B^{t+k}$. Assume $x^{t-1}\neq x^{t}$ and $x^{t+k}\neq x^{t+k+1}$. We have
$c^{T}x^{t+k+1}-c^{T}x^{\star}\leq\big(1-\frac{\delta}{(n-m)\Delta}\big)(c^{T}x^{t+k}-c^{T}x^{\star})\,.$
Proof
The optimality gap for $x^{t+k}$ can be bounded as follows:
[…]
where we used Lemma 2(a) for the first inequality, and (2) for the third equality. The last equality follows from $y^{t+k}_{i}\geq 0,i\in N^{t+k}$ which is implied by feasibility of $y^{t+k}$ at $x^{t+k}$. The last inequality follows from $\max_{\{i\in N^{t+k}\mid y^{t+k}_{i}>0\}}-\overline{c}_{N^{t+k},i}=\Delta^{t+k}_{\overline{c}}$, the fact that $|N^{t+k}|=n-m$, and because $y^{t+k}_{i}=y^{t}_{i}=x^{\star}_{i}-x^{t}_{i}\leq\Delta$ holds for any $i\in N^{t+k}\cap\,\mathrm{supp}(y^{t+k})$, using Lemma 2(b).
Let $x^{t+k}_{f}$ denote the entering variable at $(t+k)\textsuperscript{th}$ iteration. Note that $x^{t+k+1}_{f}\neq 0$ since $x^{t+k}\neq x^{t+k+1}$. Then,
$\begin{array}[]{rl}c^{T}x^{t+k}-c^{T}x^{t+k+1}&=\overline{c}_{N^{t+k}}^{T}(x^{t+k}-x^{t+k+1})_{N^{t+k}}\\ &=-\overline{c}_{N^{t+k},f}x^{t+k+1}_{f}\\ &\geq\Delta^{t+k}_{\overline{c}}\delta\\ &\geq\frac{\delta}{(n-m)\Delta}c^{T}(x^{t+k}-x^{\star})\end{array}$
hold, where we used (2) for the first equality and (6) for the last inequality. Rearranging the terms yields the lemma statement.∎
Now, combining Lemma 1 and Lemma 3 allows to conclude that the simplex algorithm with the antistalling pivot rule described in Section 2 encounters at most $(n-m)\lceil\frac{(n-m)\Delta}{\delta}\ln\big(m\frac{\Delta}{\delta}\big)\rceil$ distinct basic feasible solutions. Since Theorem 2.1 bounds the number of consecutive non-degenerate steps by $\min\{n-m-1,m-1\}$ (again, due to the choice of the improving feasible direction), combining it with the latter result yields a bound on the total number of pivots required to reach an optimal vertex. However, this bound does not take into account the number of (degenerate) pivots that might have to be performed at an optimal vertex to encounter a basis satisfying the optimality criterion. This is due to the fact that the antistalling pivot rule requires an improving feasible direction, which we do not have at an optimal vertex. Hence, we have to handle this case separately.
Lemma 4
Let $B$ and $B^{\star}$ be a non-optimal and an optimal basis, respectively, both associated with an optimal basic feasible solution $x^{\star}$ of (1). We can obtain $B^{\star}$ from $B$ by performing at most nmn-m degenerate simplex pivots.
Proof
Since $B$ is not optimal, there exists $f\in N$ with $\overline{c}_{N,f}=c^{T}z<0$ with $z$ as in (3). Observe that there exists $i\in B\cap N^{\star}$ with $\overline{A}_{if}>0$ and $x_{i}=0$: otherwise, all coordinates of $z_{N^{\star}}$ would be non-negative and hence $z\in C_{N^{\star}}:=\{x\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{n}\mid Ax=0,x_{N^{\star}}\geq 0\}$. The latter however contradicts the fact that all extreme rays of the above cone $C_{N^{\star}}$ have non-negative scalar products $\overline{c}_{N^{\star}}$ with $c$ due to optimality of $B^{\star}$. Hence one could perform a simplex pivot on $B$ with entering variable $f$ and leaving variable $i$. Let $B^{\prime}:=\big(B\setminus\{i\}\big)\cup\{f\}$ and $N^{\prime}:=\big(N\setminus\{f\}\big)\cup\{i\}$. Note that $\overline{c}_{N^{\prime},i}\geq 0$ and $i\in N^{\prime}\cap N^{\star}$. Either $B^{\prime}$ is a optimal basis and we stop, or there exists $j\in N^{\prime}\setminus\{i\}$ with $\overline{c}_{N^{\prime},j}<0$. In the latter case, however, we can enforce the constraint $x_{i}=0$ by removing the variable $x_{i}$ together with the corresponding column of $A$ and entry $c_{i}$ of $c$ from (1). By doing so we obtain a new LP with the number of variables smaller by one that has $B^{\prime}$ and $B^{\star}$ as a non-optimal and an optimal basis, respectively, since $\overline{c}_{N^{\prime}\setminus\{i\},j}<0$ and $\overline{c}_{N^{\star}\setminus\{i\}}\geq 0$. We can set $B=B^{\prime}$ and repeat the process. Since at each iteration a variable with index from $N^{\star}$ gets removed, after at most nmn-m iterations $B=B^{\star}$. ∎
Since the number of degenerate pivots at an optimal vertex is bounded by nmn-m by the above lemma, we can state the following result.
Theorem 3.1
Given an LP (1) and an initial feasible basis, there exists a pivot rule that makes the simplex algorithm reach an optimal basis in at most
$\min\{n-m,m\}(n-m)\lceil\frac{(n-m)\Delta}{\delta}\ln\big(m\frac{\Delta}{\delta}\big)\rceil$
simplex pivots.
[…]
Corollary 1
For any basic feasible solution of an LP (1) with integral $A$ and $b$, there exists a sequence of at most
[…]
simplex pivots leading to an optimal basis.
3.1 Application to combinatorial LPs
We conclude this section by observing that, using the latter corollary, one can prove the existence of short sequences of simplex pivots (that is, of length strongly-polynomial in $n,m$) between any two extreme points of several combinatorial LPs, that is, LPs modeling the set of feasible solutions of famous combinatorial optimization problems. We report a few examples below.
(a) LPs modeling matching/vertex-cover/edge-cover/stable-set problems in bipartite graphs.
For matchings, the LP maximizes a linear function over a set of constraint of the form $\{A^{\prime}x\leq 1,x\geq 0\}$, where the coefficient matrix $A^{\prime}$ is the node-edge incidence matrix of an undirected bipartite graph. After putting the LP in standard equality form by adding slack variables, we get constraints of the form $\{Ax=1,x\geq 0\}$, where $A$ is a totally unimodular matrix (and so $\Delta_{A}=1$). The result then follows from Corollary 1. The same holds for vertex-cover (minimizing a linear function over constraints of the form $\{{A^{\prime}}^{T}x\geq 1,x\geq 0\}$), edge-cover (minimizing a linear function over constraints of the form $\{{A^{\prime}}x\geq 1,x\geq 0\}$), stable-set (maximizing a linear function over constraints of the form $\{{A^{\prime}}^{T}x\leq 1,x\geq 0\}$).
(b) LPs modeling optimization over the fractional matching/vertex-cover/edge-cover/stable-set polytopes.
These correspond to the natural LP relaxations of the problems discussed in (a), for general graphs. The same LPs as in (a), in non-bipartite graphs, have a constraint matrix $A^{\prime}$ (resp. ${A^{\prime}}^{T}$) that is not totally unimodular. However, the set of constraints defines half-integral polytopes (see [2, 31]). Note that, after putting the LPs in standard equality form, the slack variables can be loosely bounded from above by the number of variables $n$. Hence, $\Delta\leq n$ and $\delta\geq\frac{1}{2}$, and the result follows from Theorem 3.1.
(c) LP for the stable marriage problem.
The classical stable marriage problem is defined on a bipartite graph where each node has a strict preference order over the set of its neighbours. One looks for a matching that does not contain any blocking pair, that is, a pair of nodes that mutually prefer each other with respect to their matched neighbour. There is an (exact) LP formulation for the problem [36, 45], that has constraints of the form $\{A^{\prime}x\leq 1,B^{\prime}x\geq 1,x\geq 0\}$. Here $A^{\prime}$ is again the node-edge incidence matrix of an undirected bipartite graph, while $B^{\prime}$ is a matrix that stems from imposing an additional constraint for each edge $\{uv\}$, that essentially prevents $\{uv\}$ from being a blocking pair:
$x_{uv}+\sum_{w:w>_{u}v}x_{uw}+\sum_{w:w>_{v}u}x_{vw}\geq 1$
In the above expression, $w>_{u}v$ (resp. $w>_{v}u$) means that $u$ prefers $w$ over $v$ (resp. $v$ prefers $w$ over $u$). After putting the LP in standard equality form, the slack variables can be bounded by $1$. Hence, $\Delta=\delta=1$, and the result follows from Theorem 3.1.
(d) LPs modeling various flow problems (such as max flows, min cost flows, flow circulations) with unit (or bounded) capacity values.
Flow problems in capacitated graphs are modeled using LPs with constraints of the form $\{{A^{\prime}}x=b,l\leq x\leq u\}$ where the constraint matrix $A^{\prime}$ is here a node-arc incidence matrix of a directed graph, which is totally unimodular (see [38]). The right-hand-side vectors ($b,l,u$) can be bounded in terms of the total capacity values. Therefore, assuming these are bounded integers, one can rely on Corollary 1 to get the result, similarly to (a).
One can compute the corresponding pivot sequence for the problems mentioned in (a)-(d) by running the simplex with the antistalling pivot rule. However, one needs to first solve the LP and obtain an optimal basic solution in order to use it to guide the antistalling pivot rule.
We remark that the existence of a strongly polynomial sequence of pivots for general min-cost flows (hence for problems in (d)), was known already, as it follows from the fundamental work of Orlin [33], that applies even in the case of large capacity values. In comparison, our bound is weaker as it requires the capacity values to be bounded.
Instead, despite being known to be solvable in strongly polynomial time, to the best of our knowledge the problems in (a)-(c) were not previously known to admit strongly polynomial bounds on the number of simplex pivots for their natural LP formulations.
4 Computational experiments
In this section we report some computational experiments that we conducted to evaluate the performance of the antistalling pivot rule when applied to actual linear programs. The code is now available at https://github.com/kirill-kukharenko/antistalling.
For the implementation, we used the python package CyLP [44] which wraps COIN-OR’s CLP solver [10] and provides tools for implementing a preferred pivot rule in python to substitute CLP’s built-in ones. We used two different testsets of diverse dimensions and sparsity. The first one is the benchmark Netlib LP dataset [32], containing 108 LPs. The second one contains the root LPs of the benchmark set of MIPLIB 2017 [17], and is composed of 240 instances. All our experiments were conducted on a laptop with Intel Core i7-13700H 2.40GHz CPU and 32GB RAM.
For each of the aforementioned LPs, we ran the simplex algorithm with our antistalling pivot rule, and in addition with other 6 well-known pivot rules (which [44] provides implementations for). These are:
  • Dantzig’s rule: the entering variable is the one with the most negative reduced cost;
[…]
  • LIFO: the entering variable is the one that minimizes the number of iterations that have past since the variable left the basis;
  • Most Frequent: the entering variable is the one that maximizes the number of times it was previously chosen as the entering variable;
  • Bland’s rule: the entering variable is the one with the lowest index.
  • Positive edge rule: the entering variable is selected with a heuristic that tries to identify non-degenerate pivots with high probability [34].
We tracked the number of pivots required by each of the above rules to solve the LPs to optimality (we did not perturb the instances). For our antistalling pivot rule, we guided it using the feasible direction $y=x^{\star}-x$ at any non-optimal vertex $x$ of the LP, for a pre-computed optimal basic solution $x^{*}$ (in particular, the one found by Dantzig’s pivot rule). Intuitively, this represents the best direction that can possibly guide our rule, as it is the direction leading immediately to an optimal solution. Our objective with the computational experiments is to see whether this choice translates into few pivots in practice, since $y$ might change during the degenerate steps and therefore in reality we do not have control on the actual edge-direction we end up moving along.
We remark that we are here interested in the number of pivots required by various pivot rules to solve LPs to optimality, and not in the computational time needed to do so. In fact, the CyLP package connects to the solver through a relatively slow Python interface, in contrast to implementing the rules directly into a solver. Therefore, the efficiency of the pivot rule implementations, and hence the number of problems solved by each rule within the time/memory limit, can be heavily affected by this overhead. For a fair comparison, in this section we then only included the instances solved to optimality by all respective pivot rules. However, for completeness, detailed tables with all the results of the computational experiments, for both Netlib and MIPLIB testsets, can be found in Appendices 0.A and 0.B, respectively.
Our first report deals with 84 out of 108 Netlib problems, for which each of the aforementioned 7 pivot rules was able to find an optimal solution within the time limit of 30 minutes and the memory limit of 32GB RAM. For the results, see Figure 2. Similarly, for MIPLIB this restriction leaves us with 117 out of 240 problems. See Figure 3 for the results.
Figure 2: Netlib dataset results. Data are presented using standard box plots, which display the distribution of a dataset based on its quartiles. Specifically, the box extends from the first quartile (Q1) to the third quartile (Q3), representing the middle $50\%$ of the data. The thin vertical lines (whiskers) extend from the box to the most extreme data points lying within 1.5 times the interquartile range (Q3-Q1). Points beyond the whiskers are shown individually as outliers (fliers). Orange horizontal lines denote the medians: 551, 659, 648, 739, 2027, 1890, and 2049, respectively.
Figure 3: MIPLIB dataset results. Orange lines denote the medians: 1376, 1633, 1493, 2335, 7374, 7829, and 8641, respectively.
The computational experiments show that the antistalling pivot rule, guided by a known optimal solution, actually performs quite well, especially on the Netlib instances. Most often it manages to find a relatively short sequence of pivots to an optimal basis compared to the other pivot rules. We highlight for example the problems 25fv47 from Netlib and istanbul-no-cutoff.mps from MIPLIB, where the antistalling pivot rule required less than one fourth of the number of iterations of any other rule. On the other hand, there are instances, e.g., fit2p from Netlib and ic97_potential.mps from MIPLIB, where the antistalling pivot rule actually showed the worst performance compared to all its competitors.
Overall, the results seem to confirm that guidance by an optimal solution indeed often translates into a small number of pivots, motivating further theoretical studies on how such a direction could be approximately followed.
5 Final Remarks
We would like to conclude the paper with some remarks and potential research directions that originate from our findings.
The results of this work show the existence of a short sequence of degenerate simplex pivots that leads out of degeneracy. It would be interesting to show that similar results can be achieved through perturbations. Informally, perturbing the right-hand-side of the constraints has the effect of “splitting” a degenerate vertex into a set of (possibly exponentially many) distinct vertices, each associated to a distinct basis. It is legitimate to ask whether every new vertex allows for a short monotone path (where monotonicity is with respect to the linear objective function of the LP), which leads to a vertex out of this set. Note that the existence of the short degeneracy-escaping sequence of simplex pivots showed in this work, generally does not answer the latter question since some bases used in the sequence might become infeasible after perturbation.
Next, we would like to connect the above question to the recent work [23]. Given a basis $B$ defining a degenerate vertex $v$, one could imagine to separate $v$ from all its neighbouring vertices by a hyperplane: this way, one would obtain a pyramid (with $v$ at its apex). Suppose this pyramid is full-dimensional and is described by some $Ax\leq b$: then, the above question is equivalent to perturbing the side facets of the pyramid to allow for a short path from the new apex defined by the given basis $B$, to any of the vertices in the bottom of the pyramid. This can be done relying on the techniques used for the construction of a so-called rock extension [23] (when viewing the base of the pyramid as the d1d-1-dimensional original polytope, and the pyramid as its $d$-dimensional extension). Such rock extension has been proven to have the property that the apex is connected to any bottom vertex by a path of at most linear length. However these paths are not necessarily monotone for a given linear objective function, and moreover, the construction of the rock extension requires not only perturbing the right-hand-side $b$ but the left-hand-side $A$ as well.
Another future research direction, that seems to be supported also by the performed computational experiments, would be to analyze the relation between the improving feasible direction $y$ that is required by our antistalling pivot rule, and the actual edge-direction $z$ along which one ends up moving after all the degenerate pivots. It would be very interesting to identify some conditions which ensure that $z$ is a good approximation of $y$, e.g., in terms of objective function’s improvement.
Furthermore, as it was mentioned in Section 2, in order for our bound to hold, the only requirement we need to impose when performing a pivot is that the entering variable lies in the support of $y$ (besides of course having negative reduced cost). In order to obtain the results from Section 3, where both degenerate and non-degenerate pivots have to be taken into account, we chose the entering variable to be the one with the most negative reduced cost, i.e., we used Dantzig’s rule just restricted to the coordinates in $\mathrm{supp}(y)$. Of course, one could think of analyzing the performance of the antistalling pivot rule with other choices of entering coordinates from $\mathrm{supp}(y)$, e.g., according to steepest edge or shadow vertex. Variations of these latter two rules, in particular, have been shown to play a key role in the context of 0/1 polytopes [4].
Acknowledgements.
The authors are grateful to Volker Kaibel for helpful discussions on the geometric interpretation of the antistalling pivot rule. We thank Alex Black and anonymous reviewers for their helpful comments. The authors express gratitude to the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) for supporting the first author within 314838170, GRK 2297 MathCoRe and providing travel funds for visiting the second author in Milan. The second author is supported by the NWO VIDI grant VI.Vidi.193.087.
Appendix 0.A Detailed Netlib results
Table 2: Netlib LPs. * signifies the instance couldn’t be solved by the respective pivot rule within the time and memory limit.
problem Dantzig PositiveEdge LIFO MostFrequent SteepestEdge Bland Antistalling
25fv47.mps 8617 7892 9934 10016 9530 9870 1468
80bau3b.mps 19858 20774 37525 38130 17628 37965 4119
adlittle.mps 111 125 197 316 98 326 56
afiro.mps 16 17 33 30 17 30 16
agg.mps 130 134 300 228 228 233 128
agg2.mps 216 188 210 194 258 200 228
agg3.mps 239 166 317 236 276 249 136
bandm.mps 674 770 2522 2762 1283 2694 674
beaconfd.mps 104 102 117 117 195 117 89
blend.mps 92 86 223 299 122 253 82
bnl1.mps 3290 2228 6295 6408 1883 6412 684
bnl2.mps 12115 * 19314 * * * 1524
boeing1.mps 675 695 2485 2570 842 2726 401
boeing2.mps 174 163 518 550 300 568 230
bore3d.mps 198 199 449 496 291 424 152
brandy.mps 353 327 1595 1701 1060 1620 182
capri.mps 506 479 561 1148 416 983 334
cre-a.mps 5828 3202 19458 25810 14105 26623 1554
cre-b.mps 326739 31408 46413 295815 * 322064 10028
cre-c.mps 6712 3504 20850 24339 17947 23915 1483
cre-d.mps 316546 22426 73494 265573 * 281915 38690
cycle.mps 5383 5152 * * * * 4621
czprob.mps 1905 2557 14585 14724 1373 14899 886
d2q06c.mps 29315 29346 30883 * * 31571 5661
d6cube.mps 22028 21286 21790 22154 22061 21891 21569
degen2.mps 1891 1342 3480 3431 729 3564 627
degen3.mps 10713 5816 12145 12312 2839 11822 4241
dfl001.mps 104120 99785 104924 105564 * 97847 73090
e226.mps 718 561 1790 1820 767 1806 216
etamacro.mps 731 609 2652 * * * 995
fffff800.mps 868 671 4169 4922 900 4943 1216
finnis.mps 611 488 810 1385 523 1163 347
fit1d.mps 2636 2369 3301 3290 2532 3377 3254
fit1p.mps 2193 1888 $56$ 7214 1340 7246 4612
fit2d.mps 29994 31236 31841 31923 * 31880 9417
fit2p.mps 30716 33042 60330 61925 11275 58919 69711
forplan.mps 243 199 * 2065 * 2055 136
ganges.mps 1628 1834 2237 2521 1801 2217 1728
gfrd-pnc.mps 1238 1030 1501 1626 1062 1813 801
greenbea.mps 28646 20538 * * * 33255 26251
greenbeb.mps 30340 13041 33436 35418 * * 30644
grow15.mps 725 647 2964 2217 2455 3016 1524
grow22.mps 1038 904 4612 3410 5427 5263 2909
grow7.mps 306 260 1157 702 838 941 469
israel.mps 331 285 453 747 431 887 123
kb2.mps 61 65 167 127 70 128 55
ken-07.mps 4622 3825 11384 18902 3748 18336 3929
ken-11.mps 31812 29710 113471 114524 * 114656 40987
ken-13.mps 222253 91988 228878 230157 * 229959 *
ken-18.mps * * * * * * *
lotfi.mps 207 287 1573 1558 295 1653 121
maros.mps 3922 2559 8837 9335 3880 8738 4375
maros-r7.mps 4144 4280 40649 44360 4176 40737 4144
modszk1.mps 1125 1082 4478 7128 1327 7132 672
nesm.mps 4392 4392 * 11946 * * 1508
osa-07.mps 600 1293 2402 2205 750 4504 1090
osa-14.mps 1400 2510 5553 4787 1500 8807 806
osa-30.mps 2617 4925 11199 14545 * 30277 1551
osa-60.mps 5234 8905 22634 149478 * 78809 2966
pds-02.mps 1776 1668 10942 16084 2099 32573 19604
pds-06.mps 28565 6813 39989 88880 8904 122043 8402
pds-10.mps 122401 13140 65131 154595 * 208463 18440
perold.mps 6339 5416 7558 * * 7697 1792
pilot.ja.mps 5565 * 10530 * * * 1354
pilot.mps * * * * * * *
pilot.we.mps 8145 3862 14524 14493 13676 14690 9037
pilot4.mps 1326 1636 4895 * * * 1409
pilot87.mps * * * * * * *
pilotnov.mps 1812 2582 12539 12187 12088 12530 907
recipe.mps 48 48 64 64 48 64 48
sc105.mps 103 102 148 120 102 119 108
sc205.mps 231 232 294 252 234 228 224
sc50a.mps 49 45 56 53 49 53 47
sc50b.mps 49 50 51 50 49 50 50
scagr25.mps 653 834 3311 3042 695 3081 771
scagr7.mps 140 186 501 361 190 401 154
scfxm1.mps 457 530 1826 1960 613 1694 537
scfxm2.mps 993 1053 4854 4990 1380 5010 2380
scfxm3.mps 1534 1575 7462 7499 2015 7519 1928
scorpion.mps 364 366 782 444 363 450 382
scrs8.mps 667 809 5386 5334 555 5474 319
scsd1.mps 272 237 2228 2753 180 2697 202
scsd6.mps 620 517 4913 4845 602 4964 237
scsd8.mps 1918 1307 10620 10652 1719 10563 2037
sctap1.mps 317 356 718 822 368 1156 191
sctap2.mps 1225 1270 1095 1352 1533 2077 658
sctap3.mps 1907 1756 1482 1626 2231 1586 944
seba.mps 560 649 508 510 876 507 976
share1b.mps 247 290 1156 1148 211 947 302
share2b.mps 157 131 245 368 156 408 216
shell.mps 733 744 1216 1268 680 1406 779
ship04l.mps 323 343 1649 1637 368 2022 366
ship04s.mps 313 328 1028 1018 346 1159 363
ship08l.mps 598 628 2469 2534 588 3413 540
ship08s.mps 566 552 1328 1335 551 1628 496
ship12l.mps $11$ 1017 6077 6234 988 8259 1152
ship12s.mps 951 914 3005 3558 919 2959 881
sierra.mps 1109 1134 970 1029 1864 1020 563
stair.mps 489 507 2884 800 503 728 1842
standata.mps 69 67 118 123 70 127 63
standgub.mps 69 67 118 123 70 127 80
standmps.mps 333 379 470 829 342 999 208
stocfor1.mps 79 79 241 268 79 247 79
stocfor2.mps 2001 2212 13526 13608 3591 13492 1774
tuff.mps 1655 720 2965 2991 701 2945 1011
vtp.base.mps 230 140 381 572 175 513 160
wood1p.mps 665 404 9584 9337 3282 9208 381
woodw.mps 2401 1877 * 32532 * 31839 2491
Table 2: Netlib
Appendix 0.B Detailed MIPLIB results
Table 3: MIPLIB root LPs. * signifies that the instance couldn’t be solved by the respective pivot rule within the time and memory limit.
problem Dantzig PositiveEdge LIFO MostFrequent SteepestEdge Blands Antistalling
30n20b8.mps 354 413 29832 51208 8610 58208 294
50v-10.mps 282 272 1114 1037 2729 1027 332
CMS750_4.mps 16295 13901 96361 100745 * 101308 41850
academictimetablesmall.mps 40401 26544 171784 176255 * * 8027
air05.mps 18529 17937 24638 25176 2531 25058 19200
app1-1.mps 2602 2488 11247 22824 7809 22766 3886
app1-2.mps 26462 25375 * * * * *
assign1-5-8.mps 314 334 1235 1199 496 1149 239
atlanta-ip.mps * 144325 * * * * *
b1c1s1.mps 8515 7124 24663 27554 13659 27456 2649
bab2.mps * * * * * * *
bab6.mps 88009 57985 * * * * *
beasleyC3.mps $13$ 1091 5691 3444 2403 2813 1430
binkar10_1.mps 1519 1493 6279 9385 1714 10168 1376
blp-ar98.mps 750 806 10794 14173 2897 52747 244
blp-ic98.mps 383 353 10432 15684 2125 43764 90
bnatt400.mps 2098 1085 5055 24275 1128 18792 2144
bnatt500.mps 3257 $16$ 6261 34694 1326 33572 6050
bppc4-08.mps 267 307 5373 4064 3543 3194 203
brazil3.mps 224690 143249 228236 * * * *
buildingenergy.mps * * * * * * *
cbs-cta.mps 3915 4629 44014 19417 * 19132 3018
chromaticindex1024-7.mps 112861 * * * * * *
chromaticindex512-7.mps 53453 * * * * * 24224
cmflsp50-24-8-8.mps 62624 29725 66646 66879 * 66483 46395
co-100.mps 1810 1432 27194 155300 * 156246 79949
cod105.mps 5043 10413 10806 * 2350 * 3517
comp07-2idx.mps 36630 19996 120086 140436 * 140131 124969
comp21-2idx.mps 17908 11221 77106 90612 3935 99479 83129
cost266-UUE.mps 1612 1782 3310 3582 4375 3626 1827
cryptanalysiskb128n5obj14.mps 83012 94390 * 125246 * * *
cryptanalysiskb128n5obj16.mps 83670 91007 * 142608 * * *
csched007.mps 1440 1902 7905 8245 1777 7794 603
csched008.mps 1012 1024 6211 6404 1265 6246 1647
cvs16r128-89.mps 28014 28833 26663 25872 * 28176 27912
dano3_3.mps 50506 * * * * * *
dano3_5.mps 50506 * * * * * *
decomp2.mps 2644 2574 7315 7555 3674 12070 5481
drayage-100-23.mps 4423 2050 15077 27251 2049 38281 6012
drayage-25-23.mps 4599 2189 14278 30007 1812 24166 6277
dws008-01.mps 313 246 1261 527 430 478 358
eil33-2.mps 233 241 1600 2129 151 2469 603
eilA101-2.mps 3325 3514 36453 79566 856 * 56899
enlight_hard.mps 29 29 29 29 29 29 29
ex10.mps * 179784 * * * * *
ex9.mps * 84528 * * * * *
exp-1-500-5-5.mps 602 1031 639 639 1311 639 457
fast0507.mps 27383 24272 196331 197020 * 197160 22921
fastxgemm-n2r6s0t2.mps 5691 3666 17063 22611 2964 19282 5782
fhnw-binpack4-4.mps 665 667 960 625 1230 615 994
fhnw-binpack4-48.mps 9521 7946 8830 4430 19723 4406 7236
fiball.mps 6688 8910 12339 79670 * 73548 4025
gen-ip002.mps 71 71 147 90 86 91 32
gen-ip054.mps 61 61 152 104 49 105 30
germanrr.mps 3879 4119 * 56995 * * 9180
gfd-schedulen180f7d50m30k18.mps * * * * * * *
glass-sc.mps 1589 2876 20767 21471 499 22452 610
glass4.mps 93 86 304 235 116 187 73
gmu-35-40.mps 1651 765 2876 1947 2183 2582 1494
gmu-35-50.mps 1955 1041 3785 3745 2992 4648 2405
graph20-20-1rand.mps 3843 2835 24081 * 2517 12187 4597
graphdraw-domain.mps 268 309 622 304 386 292 168
h80x6320d.mps 364 263 10938 8128 6955 8835 499
highschool1-aigio.mps * * * * * * *
hypothyroid-k1.mps 10923 11179 26362 15838 7999 24741 4307
ic97_potential.mps 634 598 752 514 636 519 1101
icir97_tension.mps 3186 2145 3432 4216 3587 3547 3383
irish-electricity.mps * * * * * * *
irp.mps 225 287 * 4075 139 23929 191
istanbul-no-cutoff.mps 11975 11678 82444 86299 * 80471 2781
k1mushroom.mps 81426 16692 18513 5144 * 5400 *
lectsched-5-obj.mps 154064 * 19026 65039 * 64952 45012
leo1.mps 1906 1042 10858 22634 10125 22835 275
leo2.mps 4766 2269 17501 36383 12522 36488 193
lotsize.mps 1918 1785 3498 3567 3148 3599 2010
mad.mps 37 87 97 154 52 226 40
map10.mps 91211 * * * * * *
map16715-04.mps * * * * * * *
markshare2.mps 39 39 97 110 62 107 39
markshare_4_0.mps 20 20 39 42 31 37 19
mas74.mps 129 129 516 566 307 563 23
mas76.mps 79 79 414 500 110 528 37
mc11.mps 1677 1894 3430 2643 4578 2592 1327
mcsched.mps 4708 5116 11932 11842 10099 11784 5820
mik-250-20-75-4.mps 75 75 75 75 75 75 75
milo-v12-6-r2-40-1.mps 3739 5212 26261 15404 4622 9373 2451
momentum1.mps 12355 8999 * * * * 11537
mushroom-best.mps * 57526 10307 46970 1123 * *
mzzv11.mps 39840 13789 73458 88323 * 106101 3987
mzzv42z.mps 12448 9094 62218 56988 11379 80630 3614
n2seq36q.mps 4060 2892 77923 81065 3048 79541 4388
n3div36.mps 143 150 2276 5299 295 8577 37
n5-3.mps 1396 995 11634 11481 2012 11550 890
neos-631710.mps 31810 17901 * * * * *
neos-662469.mps 30106 23253 61914 63061 * 62946 50345
neos-787933.mps 2163 2246 * 9279 * 8700 2164
neos-827175.mps 171542 * * 165678 * 165157 *
neos-848589.mps 2135 * * 13758 * * 2114
neos-860300.mps 1155 893 7698 7829 1321 7938 224
neos-873061.mps 110262 * * * * * 33
neos-911970.mps 129 133 3183 3363 214 3447 92
neos-933966.mps 158717 167025 169776 155596 * 158554 109968
neos-950242.mps 6928 7299 64874 80748 2590 5921 2630
neos-957323.mps 28302 12481 * * * * *
neos-960392.mps 9058 10795 $203$ 205651 * 206419 8805
neos17.mps 936 767 3430 1605 570 1379 700
neos5.mps 105 107 193 122 99 270 7
neos8.mps 712 429 4581 6209 553 9684 1124
neos859080.mps 4 4 93 91 83 90 4
neos-1122047.mps 8685 8825 * * * * 9201
neos-1171448.mps 79780 71487 59813 6001 * 5843 18055
neos-1171737.mps 25462 21642 21740 2542 22777 2630 22927
neos-1354092.mps 90484 88911 67471 67230 * 61909 15903
neos-1445765.mps 6122 6555 28656 31982 10084 46889 5772
neos-1456979.mps 35773 8662 30198 35218 2142 35123 3378
neos-1582420.mps 3670 2964 47584 63690 6191 64109 4293
neos-2075418-temuka.mps * * * * * * *
neos-2657525-crna.mps 360 200 627 699 305 678 292
neos-2746589-doon.mps 120878 36276 * 271882 * * 14163
neos-2978193-inde.mps 928 2709 37226 64529 2335 64472 833
neos-2987310-joes.mps 14222 13281 23023 25572 * 26282 15458
neos-3004026-krka.mps 6660 6660 6305 6305 * 6305 11610
neos-3024952-loue.mps 3645 3863 13076 $45$ 8948 5767 3073
neos-3046615-murg.mps 169 169 235 235 455 235 169
neos-3083819-nubu.mps 2678 2409 8224 11544 2240 15352 1681
neos-3216931-puriri.mps * 34396 45197 45797 9782 50236 *
neos-3381206-awhea.mps 1584 1774 2196 2312 1978 2227 2002
neos-3402294-bobin.mps 1066 1023 3374 * 461 * 1134
neos-3402454-bohle.mps 5481 3457 * * * * *
neos-3555904-turama.mps 6076 1607 * * * * 4329
neos-3627168-kasai.mps 1633 1900 4127 3905 5382 5776 1217
neos-3656078-kumeu.mps 20525 17539 113007 110161 * 109498 9260
neos-3754480-nidda.mps 291 430 2202 974 635 1035 231
neos-3988577-wolgan.mps 40469 24658 * * * * *
neos-4300652-rahue.mps 85331 17034 48186 * * * 26589
neos-4338804-snowy.mps 683 692 441 441 792 441 1681
neos-4387871-tavua.mps 2292 2471 9202 6459 3000 26601 3427
neos-4413714-turia.mps 27839 10821 * * * * 16425
neos-4532248-waihi.mps 3191 982 * * * * 7981
neos-4647030-tutaki.mps 11605 12066 18514 18452 * 18167 *
neos-4722843-widden.mps 14983 10825 * 30597 * 37043 *
neos-4738912-atrato.mps 4027 5190 26204 26796 27602 26215 5533
neos-4763324-toguru.mps 302123 71563 * * * * *
neos-4954672-berkel.mps 962 965 3574 3792 3055 4385 670
neos-5049753-cuanza.mps * * * * * * *
neos-5052403-cygnet.mps * * * * * * *
neos-5093327-huahum.mps 19617 20575 * * * * *
neos-5104907-jarama.mps * * * * * * *
neos-5107597-kakapo.mps 6191 $69$ 14446 11175 * 10175 8147
neos-5114902-kasavu.mps * * * * * * *
neos-5188808-nattai.mps 26125 12571 24862 46541 * * 9429
neos-5195221-niemur.mps 43996 14654 21534 * * * 13985
net12.mps 3929 5043 29848 92171 4716 92870 1160
netdiversion.mps * * * * * * *
nexp-150-20-8-5.mps 661 650 2707 5219 11536 6470 845
ns1116954.mps 5600 12858 441155 24617 * 81870 19024
ns1208400.mps 3263 2398 25447 25859 2281 26724 $45$
ns1644855.mps * * * * * * *
ns1760995.mps * 31060 * * * * *
ns1830653.mps 18120 16096 18666 19139 4744 19311 11198
ns1952667.mps 68 68 11145 18491 87 24957 68
nu25-pr12.mps 1850 1642 6699 26414 3099 26281 6165
nursesched-medium-hint03.mps 156192 36390 160684 169581 * 168361 35869
nursesched-sprint02.mps 16279 8259 47487 46400 3100 46185 46907
nw04.mps 438 313 12328 56821 130 * 94
opm2-z10-s4.mps * * 28984 * * * *
p200x1188c.mps 578 579 1603 1061 1177 1018 1244
peg-solitaire-a3.mps 11502 10055 31045 30794 6902 30915 19201
pg.mps 1566 1567 3628 3730 8714 4477 1552
pg5_34.mps 1684 1999 8040 8664 3631 8641 3022
physiciansched3-3.mps * * * * * * *
physiciansched6-2.mps * * * * * * *
piperout-08.mps 10035 8071 77316 78677 13883 79658 10936
piperout-27.mps 13212 10491 93333 97329 * 98047 19730
pk1.mps 56 56 234 256 229 220 56
proteindesign121hz512p9.mps 963 905 13935 19537 * * 1311
proteindesign122trx11p8.mps 606 652 11270 10568 2918 10460 713
qap10.mps 43368 * 47037 42295 * 51073 8507
radiationm18-12-05.mps 40126 50952 * * * * *
radiationm40-10-02.mps 163087 * * * * * *
rail01.mps * * * * * * *
rail02.mps * * * * * * *
rail507.mps 24907 20067 195912 196516 * 197922 23777
ran14x18-disj-8.mps 1840 1657 3587 3867 3889 3821 2611
rd-rplusc-21.mps 475 412 342 913 640 840 517
reblock115.mps 2456 5185 11934 2566 4122 2627 3220
rmatr100-p10.mps 3409 4706 35091 15813 1672 17698 501
rmatr200-p5.mps 65929 85034 * * * * 309
rocI-4-11.mps 9498 13144 46230 54614 7263 42979 20292
rocII-5-11.mps 6064 1315 3874 2342 1884 2382 1615
rococoB10-011000.mps 20776 1353 21938 21425 3977 21581 11982
rococoC10-001000.mps 3941 1511 7374 13936 5032 14324 522
roi2alpha3n4.mps 24697 10346 6264 32500 39804 10864 9391
roi5alpha10n8.mps 12459 6064 13999 * * 132274 *
roll3000.mps 1815 2155 12717 13269 1892 13997 3166
s100.mps * * * * * * *
s250r10.mps * * * * * * *
satellites2-40.mps 73896 51346 19366 213784 * 21851 2498
satellites2-60-fs.mps 47644 41964 18033 181057 * 16223 2921
savsched1.mps * * * * * * *
sct2.mps 6949 2983 27287 26895 7397 27005 3571
seymour.mps 7345 9261 22067 24986 2502 24940 15189
seymour1.mps 7345 9261 22067 24986 2502 24940 15189
sing44.mps 261483 * * * * * *
sing326.mps 220981 * * * * * *
snp-02-004-104.mps * * * * * * *
sorrell3.mps 1397 1356 1772 1369 1335 1368 1736
sp150x300d.mps 318 310 577 479 484 475 307
sp97ar.mps 3950 3573 40951 50077 2571 49995 14935
sp98ar.mps 20342 10799 39116 52886 2519 52395 1967
splice1k1.mps 35351 17482 34404 48443 * 32872 16224
square41.mps * * * * * * *
square47.mps * * * * * * *
supportcase10.mps * * * * * * *
supportcase12.mps * * * * * * *
supportcase18.mps 1787 1077 22628 27455 652 37010 499
supportcase19.mps * * * * * * *
supportcase22.mps * 2431 * * * * *
supportcase26.mps 294 313 365 409 450 409 540
supportcase33.mps 4216 3176 44224 140227 * 87128 5053
supportcase40.mps 100160 38705 * 175363 * 181878 21721
supportcase42.mps * * * * * * *
supportcase6.mps 9576 3895 * * * * 931
supportcase7.mps 154110 24378 * * * * *
swath1.mps 640 1049 19946 23793 281 23712 479
swath3.mps 640 1049 19946 23793 281 23712 433
tbfp-network.mps 42643 11572 244825 237023 * 237077 3225
thor50dday.mps 38456 * * * * * *
timtab1.mps 328 250 744 812 502 751 175
tr12-30.mps 696 1321 3304 5668 4135 5688 696
traininstance2.mps 13397 13713 21531 21904 * 32415 15256
traininstance6.mps 10270 10282 16929 15868 11367 52302 10512
trento1.mps 57604 74004 40969 40642 * 40722 34730
triptim1.mps * * * * * * *
uccase12.mps * * * * * * *
uccase9.mps 116722 * * * * * *
uct-subprob.mps 9495 7117 14877 16206 2416 16033 9294
unitcal_7.mps 150003 84117 * * * * 31193
var-smallemery-m6j6.mps 59701 59539 28200 60575 * 61423 *
wachplan.mps 4655 5722 17255 17440 2080 16360 6012
Table 3: MIPLIB